# Data Compression for Fast Online Stochastic Optimization

### v1

@@ 2 The online stochastic problem
While incorporating streaming data improves the solution quality, it also increases the computational burden, scaling with the number of data-points [40]. We thus propose a data-compression framework for fast online DRO. Our goal is to solve an adjusted DRO problem over time, where
$H^{K}_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],\quad t=1,2,\dots$ (6)
@@
and $\mathcal{P}^{K}_{t-1}$ is an adaptive ambiguity set, with adaptive radius $\varepsilon^{K}_{t-1}\geq 0$, constructed from the data $\mathcal{D}_{t-1}$ observed so far. From here onward, we refer to the solution $x^{K}_{t}$ as the online solution, and the solution $x_{t}$ from (4) as the nominal DRO solution.
@@
$\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{p}(\hat{\mathbf{P}}^{K}_{t})=\left\{\mathbf{Q}\in\mathcal{W}({S})\mid W_{p}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q})\leq\varepsilon^{K}_{t}\right\}.$ (7)
Unlike the standard DRO ambiguity set, the adaptive ambiguity set is compressed in that we allow the reference distribution to be supported on only $K\leq n_{t}$ points, to greatly reduce the computational burden. With access to the dataset $\mathcal{D}_{t}$, the reference distribution is formulated by clustering the data-points into sets $C^{k}_{t}$, $k=1,\dots,K_{t},$ where $K_{t}\leq K$. The weight $\theta^{k}_{t}$ of a cluster $C^{k}_{t}$ is defined as the proportion of total points in the cluster, and the cluster centroid $\bar{u}^{k}_{t}$ is the mean of these points. These values characterize the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$i.e.,
@@
With an adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$, we solve the online compressed DRO problem (6) through a direct reformulation approach [56, Section 2.4], given as
$\begin{array}[]{lll}\!\!H^{K}_{t}\!=\!\!\!&\underset{x\in\mathcal{X},\lambda_{t}\geq 0,s_{t},z_{t},y_{t}}{\mbox{minimize}}&\sum_{k=1}^{K_{t}}\theta^{k}_{t-1}s^{k}_{t}\\ &\mbox{subject to}&[-f_{j}]^{*}(z^{jk}_{t}-y^{jk}_{t},x)+\sigma_{{S}}(y^{jk}_{t})-(z{{}^{jk}_{t}})^{T}\bar{u}^{k}_{t}+\phi(q)\lambda_{t}\|z^{jk}_{t}/\lambda_{t}\|_{*}^{q}\\ &&\hskip 80.00012pt+\lambda_{t}\varepsilon_{t-1}^{p}\leq s^{k}_{t},\quad k=1,\dots,K_{t},\,j=1,\dots,J\\ \end{array}\!\!\!\!\!\!\!\!\!\!\!\!\!$ (8)
@@
    We compute decision $x^{K}_{t}$ by solving (8) with ambiguity set $\mathcal{P}^{K}_{t-1}$;
@@
    We update the ambiguity set to $\mathcal{P}^{K}_{t}$ with the new data point.
To assess the performance of our online solution scheme, in Section 3 we provide finite-sample and asymptotic performance guarantees of a similar form as (5), where we analyze the effects of clustering. In Section 4, we further consider the dynamic regret, which measures the performance gap between our online solution and the DRO solution in hindsight at time step $t$, both evaluated with respect to the non-compressed and updated ambiguity set $\mathcal{P}_{t}$. The notion of regret is in line with literature on online learning [3, 44, 9], where a time-discrepancy (which indicates an information-discrepancy) is introduced between the online solution and an oracle. Formally, we quantify this gap as
$R(T,K)=\frac{1}{T}\sum_{t=1}^{T}\left(\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]-\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]\right),$ (9)
where we compare the average performance of our online solution $x^{K}_{t}$ against the average performance of the oracle: the DRO solution in hindsight.
@@ 3 Clustering effect analysis
In this section, we state our main results: the effect of clustering on the finite samples and asymptotic performance guarantees for the online algorithm. These are confidence bounds, similar to (5), on the in-sample and out-of-sample performance of the online solution $x^{K}_{t}$. To provide the preliminaries, we break this section into several subsections. In Section 3.1, we state key measure concentration results. Next, in Section 3.2, we provide requirements on the online clustered distribution $\hat{{\mathbf{P}}}^{K}_{t}$; later, in Section 6, we provide specific algorithms to formulate this distribution. In Section 3.3, we then analyze this clustered distribution, including its distance from the empirical distribution $\hat{{\mathbf{P}}}_{t}$, its convergence to a limiting distribution, and its convergence rate. Finally, making use of the above information, in Section 3.4 we give the main results.
3.1 Measure concentration and nominal results
In order to obtain performance guarantees, we require the distribution ${\mathbf{P}}$ to satisfy certain assumptions. We impose the following light-tailed assumption, which is satisfied by the common class of sub-Gaussian distributions, and by bounded distributions.
Assumption 3.1 (Light-tailed distribution).
There exists an exponent $r>1$, such that $R={\mbox{\bf E}}_{\mathbf{P}}[\exp(\|u\|^{r})]=\int_{S}\exp(\|u\|^{r}){\mathbf{P}}(du)<\infty.$
This leads to the following measure concentration results.
Theorem 3.1 (Measure concentration, light-tailed [20]).
If Assumption 3.1 holds and $p=1$, for the distribution ${\mathbf{P}}$ and empirical distribution $\hat{{\mathbf{P}}}^{N}$ supported on $N$ points, we have
$\mathbf{P}^{N}(W_{1}(\mathbf{P},\hat{\mathbf{P}}^{N})\geq\varepsilon)\leq\begin{cases}c_{1}\exp(-c_{2}N\varepsilon^{\max\{d,2\}})\quad\textit{if }\varepsilon\leq 1\\ c_{1}\exp(-c_{2}N\varepsilon^{r})\quad\textit{if }\varepsilon>1,\end{cases}$
for all $\varepsilon>0$, $d\neq 2$, and $N\geq 1$. When $p\geq 2$ and $r\in(0,p)$, we have
$\mathbf{P}^{N}(W_{p}(\mathbf{P},\hat{\mathbf{P}}^{N})\geq\varepsilon)\leq\begin{cases}c_{1}\exp(-c_{2}N\varepsilon^{\max\{d,2p\}})+\exp(-c_{2}(N\varepsilon^{p})^{\frac{r-s}{p}})\\ \hfill\textit{if }\varepsilon\leq 1,\forall s\in(0,r)\\ c_{1}\exp(-c_{2}(N\varepsilon^{p})^{r/p})\quad\textit{if }\varepsilon>1,\end{cases}$
for all $\varepsilon>0$, $d\neq 4$, and $N\geq 1$. For all cases, $c_{1},c_{2}$ are positive constants depending only on $R,r,s,$ and $d$.
For the special case $p=\infty$, we point to [50, Theorem 1.1] for an analogous result, which assumes a stronger condition than Assumption 3.1. Using these relations, the radius for a Wasserstein DRO problem with $N=n_{t-1}$ points can be chosen by setting the right-hand-side to $\beta_{t-1}$, then solving for $\varepsilon_{t-1}$. See [40, Theorem 3.4] for the case $p=1$. The procedure for $p\geq 2$ follow similarly. In Section 5, we further calculate explicit radii for the specific case of bounded support $S$. In all cases, we note that the radius is inversely related to the number of data-points, i.e., the larger the number $n_{t}$, the smaller $\varepsilon_{t}$ can be. With the radius chosen in this way, we obtain the DRO performance guarantees (5) using [40, Theorem 3.5].
From the above characterization, it is evident that larger datasets are beneficial for reducing the distributional ambiguity around the nominal distribution. Indeed, for $n_{t}\rightarrow\infty$, we can choose a confidence sequence $\beta_{t}\in(0,1)$ such that $\sum_{t=1}^{\infty}\beta_{t}<\infty$ and $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ to conclude that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\hat{\mathbf{P}}_{t})\}=1$, which then implies $\lim_{t\rightarrow\infty}H_{t}\downarrow H_{\star}$ [40, Theorem 3.7]. However, the increase in the number of data-points makes the resolution of the DRO problem in (4) computationally challenging, as a constraint needs to be added for each data-point constituting the empirical distribution $\hat{\mathbf{P}}_{t}$ (see the problem formulation (8)). Clearly, for $t\rightarrow\infty$, we would end up with an infinite-dimensional problem that cannot be solved.
@@ 3.2 Clustering procedure requirements
To address the above dilemma, the key of our approach is the adaptive ambiguity set $\mathcal{P}^{K}_{t}$, which is constructed using a compressed representation of the available data, $\hat{{\mathbf{P}}}^{K}_{t}$. In this section, we formalize the requirements on generating $\hat{{\mathbf{P}}}^{K}_{t}$, and note that the decision-maker has the flexibility to use any method to generate $\hat{{\mathbf{P}}}^{K}_{t}$, provided it satisfies the given assumptions. These requirements ensure the validity of the asymptotic performance guarantees established later in this section, and can be satisfied by any partitioning of the support set $S$.
Regardless of the method, the overall goal is to find the clustering that minimizes the discrepancy between the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ and the true distribution ${\mathbf{P}}$, which is approximated by the empirical distribution $\hat{{\mathbf{P}}}_{t}$. In this section, we also outline the theoretical procedure to find the optimal clustering, and note that the goal of the online procedure is to find the best approximation of the optimal clustering, while also maintaining a low computational complexity.
@@ 3.2.1 Requirements
For each cluster $k$ and time $t$, the cluster support $S_{t}^{k}\subset S$ is defined as a region in $S$ such that any data-point $\hat{u}\in\mathcal{D}_{t}$ that falls within this region is contained in cluster $C_{t}^{k}$. We also call these clusters the clustering induced by the cluster support.
@@
  • The elements in the set $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}}$ have pairwise disjoint interiors, i.e.,
    $\mathrm{int}(S^{k}_{t})\cap\mathrm{int}(S^{k^{\prime}}_{t})=\emptyset,~{}\forall S^{k},S^{k^{\prime}}\in\mathcal{S}_{t},k\neq k^{\prime}.$
@@
  • The cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{t})>0$ for $k=1,\dots,K_{t}$ and for all $t$.
@@
These assumptions ensure that all data-points fall in a single cluster support, and could be satisfied by any partitioning of the original support set $S$. Note that until some finite time $\tau$, the data-points are allowed to switch cluster assignments, which increases the flexibility of the clustering algorithm. By Assumption 3.2, the cluster weights $\sum_{k=1}^{K_{t}}\theta^{k}_{t}=1$, and $\hat{\mathbf{P}}^{K}_{t}$ represents an approximation of the non-clustered empirical distribution $\hat{\mathbf{P}}_{t}$, supported on $K_{t}\leq n_{t}$ atoms.
@@ 3.2.2 Optimal clustering
If the distribution ${\mathbf{P}}$ is known, then for any given $K$, the optimal set of supports can be determined using the $k$-centers clustering approach described in [24, Chapter 1]. Specifically, we solve the $k$-centers problem to obtain an optimal set of centers $A^{\star}\subset S$, then construct the Voronoi diagram of $A^{\star}$. Each cluster support is defined as $S^{k}=\left\{u\in S\,\middle|\,\|u-a^{k}\|=\min_{a\in A^{\star}}\|u-a\|\right\},$ where each point $u$ is assigned to the region corresponding to its closest center $a^{k}$.
Definition 3.2 ($k$-centers problem).
The $k$-centers problem, with respect to a distribution ${\mathbf{P}}$ and orders $p\geq 1$ (left), $p=\infty$ (right) is given as
$\inf_{A\subset S,|A|\leq K}{\mbox{\bf E}}_{\mathbf{P}}\left[\min_{a\in A}\|u-a\|^{p}\right],~{}\quad\inf_{A\subset S,|A|\leq K}\sup_{u\in S}\min_{a\in A}\|u-a\|,$ (10)
@@
We also define the minimium quantization error of ${\mathbf{P}}$, and note a subsequent proposition.
Definition 3.3 (Minimum quantization error [24, Chapter 1]).
The minimum quantization error of a distribution ${\mathbf{P}}$ is the minimum Wasserstein-$p$ distance between ${\mathbf{P}}$ and any discretization of ${\mathbf{P}}$ to at most $K$ atoms, given as
$d^{K}_{\star,p}({\mathbf{P}})=\left({\mbox{\bf E}}_{\mathbf{P}}\left[\min_{a\in A^{\star}}\|u-a\|^{p}\right]\right)^{1/p},\quad d^{K}_{\star,\infty}=\sup_{u\in S}\min_{a\in A^{\star}}\|u-a\|,$
where $A^{\star}$ is the set of optimal centers.
Proposition 3.1.
[…]
Proof.
This holds since $\hat{{\mathbf{P}}}^{K}_{t}$ is a particular discretization of ${\mathbf{P}}$ to $K$ atoms. ∎
Without knowledge of ${\mathbf{P}}$ but having access to a dataset $\mathcal{D}_{t}$, the optimal clustering with respect to the empirical distribution $\hat{\mathbf{P}}_{t}$ can also be found through the above process, where ${\mathbf{P}}$ is replaced by $\hat{\mathbf{P}}_{t}$. However, the $k$-centers problem is NP-hard [29], and is therefore inefficient to solve. Furthermore, with $\hat{\mathbf{P}}_{t}$ as the reference distribution, when $t\rightarrow\infty$, the computational complexity of the $k$-centers problem increases, conflicting with our goal of decreasing computational effort. Thus, the goal of the online clustering algorithm is to find a good approximation of the $k$-centers problem, while also minimizing computational complexity. In Section 6, we give a few algorithms that satisfy the given requirements.

### v2

@@ 2.3 Our online algorithm
We propose a data-compression framework for fast online DRO. Our goal is to solve an adjusted DRO problem over time, where $K$ is the chosen number of data clusters, and
$H^{K}_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],\quad t=1,2,\dots$ (7)
@@
with $\mathcal{P}^{K}_{t}$ an adaptive ambiguity set, with adaptive radius $\varepsilon^{K}_{t}\geq 0$, constructed from the data $\mathcal{D}_{t}$ observed so far. From here onward, we refer to the solution $x^{K}_{t}$ as the online solution, and the solution $x_{t}$ from (4) as the nominal DRO solution.
@@
$\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{r}(\hat{\mathbf{P}}^{K}_{t})=\left\{\mathbf{Q}\in\Xi_{r}({S})\mid W_{r}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q})\leq\varepsilon^{K}_{t}\right\}.$ (8)
Unlike the standard DRO ambiguity set, the adaptive ambiguity set is compressed in that we allow the reference distribution to be supported on only $K\leq n_{t}$ points. With access to the dataset $\mathcal{D}_{t}$, the reference distribution is formulated by clustering the datapoints into sets $C^{k}_{t}$, $k=1,\dots,K_{t},$ where $K_{t}\leq K$. The weight $\theta^{k}_{t}$ of a cluster $C^{k}_{t}$ is defined as the proportion of total points in the cluster, and the cluster mean $\bar{u}^{k}_{t}$ is the mean of these points. These values characterize the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$i.e.,
@@
With an adaptive ambiguity set $\mathcal{P}^{K}_{t}$, we solve the online compressed DRO problem (7) through a direct reformulation approach [54, Section 2.4], given as
@@
In addition to the primal variables $x\in\mathcal{X}$, we introduce dual and auxiliary variables $\lambda_{t}\in{\bf R}$, $s^{k}_{t}\in{\bf R}$, $z^{jk}_{t}\in{\bf R}^{d}$, and $y^{jk}_{t}\in{\bf R}^{d}$, for $k=1,\dots,K_{t}$ and $j=1,\dots,J$. Here, $f_{j}]^{*}(z,x)=\sup_{u\in\mathop{\bf dom}_{u}f}z^{T}u-[-f_{j}(u,x)]$ is the conjugate of $f_{j}$, $\sigma_{S}(y)=\sup_{u\in S}z^{T}u$ is the support function of $S\subseteq\mathbf{R}^{m}$, $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|$, and $\phi(q)=(q-1)^{(q-1)}/q^{q}$ for $q>1$ [33, Theorem 8]. Note that $q$ is the conjugate number of $r$, satisfying $1/r+1/q=1$, i.e., q=r/(r1)q=r/(r-1). The above reformulation is for $r<\infty$; the case for $r=\infty$ is given in [54, Appendix B]. We remark that the computational complexity of this problem is directly correlated to the value $K_{t}$, as it controls the number of constraints required.
@@
    We compute decision $x^{K}_{t}$ by solving (9) with ambiguity set $\mathcal{P}^{K}_{t}$;
@@
    We update the ambiguity set to $\mathcal{P}^{K}_{t+1}$ with the new data point.
Remark on the re-optimization frequency
Above, we proposed to update the ambiguity set and solve (9) after each new datapoint is collected. This could be adjusted, however, depending on the frequency of data collection and the complexity of (9). If the frequency is high, or if the data arrives in batches, step 1 (and 3, if desired) of the sequential process could instead be performed once a threshold number of datapoints has been observed. This does not affect the theoretical results in Section 3.3.2, and we do so in the example in Section 7. We can also choose to solve the robust problem only after a large enough adjustment to $\hat{{\mathbf{P}}}^{K}_{t}$ has been observed: if solving (9) is time-consuming relative to data-collection, we might wish to re-optimize only if a center $\bar{u}^{k}_{t}$ or a probability $\theta^{k}_{t}$ undergoes a change beyond a predetermined threshold $\delta>0$, i.e., when $|\bar{u}^{k}_{t}-\bar{u}^{k}_{\text{prev-solve}}|\geq\delta$ for some cluster $k$. Checking for this is at most $O(K)$ at each time step. Overall, since the computational complexity of our problem can be controlled by $K$, while the complexity of nominal DRO, and even SAA, increases over time, our method enjoys much higher flexibility in choosing a re-optimization threshold.
Our performance guarantees
To assess the performance of our online solutions, in Section 3.3 we provide finite-sample and asymptotic performance guarantees similar to (5), with adjustments due to clustering. A key strength of our approach is its ability to utilize established performance guarantees for Wasserstein DRO, with minimal additional assumptions. Informally, if we choose $\varepsilon_{t}^{K}=\varepsilon_{t}$ for a nominal radius $\varepsilon_{t}$ (e.g. from Theorem 2.1 or Theorem 2.2), we obtain
${\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\rho_{t}+\psi^{K}_{t}\leq H_{t}+\rho_{t}+\gamma^{K}_{t}\right)\geq 1-\beta_{t},$ (10)
where $\psi^{K}_{t}$ and $\gamma^{K}_{t}$ are additional residual terms of order $O(\varepsilon_{t}+d(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}_{t}^{K}))$, for some distance $d(\cdot)$. In other words, the difference between our certificates and the nominal DRO certificates converges, with rate $O(\varepsilon_{t})$, to a distance between the empirical and clustered empirical distributions. This rate depends on the radius $\varepsilon_{t}$, which we have shown to scale with $O(n_{t}^{-r/d})$(dimension-dependent), or with $\tilde{O}(n_{t}^{-1/2})$ for $r\in[1,2]$ and any $d$ (dimension-free). Our residual terms can be interpreted as the price for higher computational efficiency.
@@ 3 Clustering effect analysis
In this section, we formally show the effect of clustering on the finite samples and asymptotic performance guarantees for the online algorithm. In Section 3.1, we provide requirements on the online clustered distribution $\hat{{\mathbf{P}}}^{K}_{t}$, and in Section 3.2, provide theoretical results pertaining this distribution. In Section 3.3, we state our performance guarantees.
@@ 3.1 Clustering procedure requirements
The key of our approach is the adaptive ambiguity set $\mathcal{P}^{K}_{t}$, which is constructed using a compressed representation of the available data, $\hat{{\mathbf{P}}}^{K}_{t}$. Below, we formalize the requirements on generating $\hat{{\mathbf{P}}}^{K}_{t}$, and note that the decision-maker has the flexibility to use any method to generate $\hat{{\mathbf{P}}}^{K}_{t}$, provided it satisfies the given assumptions. These requirements ensure the validity of the asymptotic performance guarantees established later in this section, and can be satisfied by any partitioning of the support set $S$. In this section, we also outline the theoretical procedure to find the optimal clustering, and note that the goal of the online clustering procedure is to find a good approximation of the optimal clustering, while also maintaining low computational complexity.
@@ 3.1.1 Requirements
For each cluster $k$ and time $t$, the cluster support $S_{t}^{k}\subset S$ is defined as a region in $S$ such that any datapoint $\hat{u}\in\mathcal{D}_{t}$ that falls within this region is contained in cluster $C_{t}^{k}$. We also call these clusters the clustering induced by the cluster support.
@@
  • The elements in the set $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}}$ have pairwise disjoint interiors, i.e., $\mathrm{int}(S^{k}_{t})\cap\mathrm{int}(S^{k^{\prime}}_{t})=\emptyset,~\forall S_{t}^{k},S_{t}^{k^{\prime}}\in\mathcal{S}_{t},k\neq k^{\prime}.$
@@
  • The cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{t})>0$ for $k=1,\dots,K_{t}$.
  • The cluster boundaries have measure 0.
@@
These assumptions ensure that all datapoints fall in a single cluster support, and could be satisfied by any partitioning of the original support set $S$. Note that until some finite time $\tau$, the datapoints are allowed to switch cluster assignments, which increases the flexibility of the clustering algorithm. By Assumption 3.1, the cluster weights $\sum_{k=1}^{K_{t}}\theta^{k}_{t}=1$, and $\hat{\mathbf{P}}^{K}_{t}$ represents an approximation of the non-clustered empirical distribution $\hat{\mathbf{P}}_{t}$, supported on $K_{t}\leq n_{t}$ atoms.
@@ 3.1.2 Optimal clustering
If the distribution $\mathbf{P}$} is known, then for any $K$, the optimal set of supports can be determined using $k$-centers clustering [20, Chapter 1].
Definition 3.2 ($k$-centers).
The $k$-centers problems, with respect to a distribution $\mathbf{Q}$ with support $Z$, and orders $p\geq 1$ (left), $p=\infty$ with compact $Z$ (right) are
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{A\subset{\bf R}^{d},|A|\leq K}{\mbox{\bf E}}_{\mathbf{Q}}\left[\min_{a\in A}\|u-a\|^{p}\right],~\quad\text{and}\quad\inf_{A\subset{\bf R}^{d},|A|\leq K}\sup_{u\in Z}\min_{a\in A}\|u-a\|,$} (11)
@@
Definition 3.3 (Minimum quantization error [20, Chapter 1]).
The minimum quantization error of a distribution $\mathbf{Q}$ with support $Z$ is the minimum Wasserstein-$p$ distance between $\mathbf{Q}$ and any discretization of $\mathbf{Q}$ to at most $K$ atoms, given as
$W^{K}_{\star,p}(\mathbf{Q})=\left({\mbox{\bf E}}_{\mathbf{Q}}\left[\min_{a\in A^{\star}}\|u-a\|^{p}\right]\right)^{1/{p}},\quad\text{and}\quad W^{K}_{\star,\infty}(\mathbf{Q})=\sup_{u\in Z}\min_{a\in A^{\star}}\|u-a\|,$
with $p\geq 1$ (left), $p=\infty$ with compact $Z$ (right), and $A^{\star}$ the set of optimal centers.
Specifically, we solve the $k$-centers problem to obtain an optimal set of centers $A^{\star}\subset{\bf R}^{d}$, then construct the Voronoi diagram of $A^{\star}$: each cluster support is defined as $S^{k}=\left\{u\in S\,\middle|\,\|u-a^{k}\|=\min_{a\in A^{\star}}\|u-a\|\right\},$ where each point $u$ is assigned to the region corresponding to its closest center $a^{k}$. This achieves the minimum quantization errori.e., $W^{K}_{\star,p}({\mathbf{P}})\leq W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}})$ for any valid $\hat{{\mathbf{P}}}^{K}_{t}$.
Without knowledge of $\mathbf{P}$} but having access to a dataset $\mathcal{D}_{t}$, the optimal clustering with respect to the empirical distribution $\hat{\mathbf{P}}_{t}$ can also be found through the above process. However, the $k$-centers problem is NP-hard [26], and is therefore inefficient to solve. Furthermore, with $\hat{\mathbf{P}}_{t}$ as the reference distribution, when $t\rightarrow\infty$, the computational complexity of the $k$-centers problem increases, conflicting with our goal of decreasing computational effort. Thus, the goal of the online clustering algorithm is to find a good approximation of the $k$-centers problem. In Section 5, we give three algorithms that satisfy the given requirements.